CookBooks Inc.’s adjusted trial balance showed total debits and total credits of $10,000 each. The company’s revenues were $6,000, its expenses were $4,000, and it declared $500 in dividends during the year. What will the debits (and credits) equal on the post-closing trial balance?
Answer #1

The debits and credits equal on the post closing trial balance will be = $5,500

Explanation: On the debit side the amount of expenses of $4,000 and dividend of $500 will be closed. Thus debit balance = 10,000 - 4000 - 500 = $5500

On the credit side revenue account will be closed and retained earnings will be added. Retained earnings = 6,000 - 4,000 - 500 = $1500. Thus balance on credit side = 10,000 - 6,000 + 1500

=$5500